# What You See Is Not Always What You Get: An Empirical Study of Code Comprehension by Large Language Models

### v2

@@ 4.4. Attack generation
4.5. Prompting
We craft the prompts with the zero-shot prompting technique. We ensure that the templates of each prompt remain consistent across all categories of perturbations and all models. Following the implementation of different attack strategies, we leverage the Chat Completions API provided by the OpenAI API to access the tested LLMs, which can be found in the replication package. The prompt template consists of the following three messages:
  1. (1)
@@ 4.5. Prompting
We note the similar question in the study of (Fang et al., 2023), which is open-ended. However, due to the different methodology design and evaluation scheme, we reword the question in a subtly different way, which only allow a one-word answer. Moreover, the code and description were taken from the LeetCode dataset. The description is thus treated as the ‘base truth’, in that it can be guaranteed due to the nature of the dataset that the correct answer to the question is always ‘Yes’.
4.6. Logprobs
To present a thorough data analysis, we used the ‘logprobs’ parameter of the Chat Completions API as the main vector. With logprobs enabled, the Chat Completions API returns the log probabilities of each output token in the response. Log probabilities of output tokens indicate the likelihood of each token occurring in the sequence (log, 2024). A higher log probability suggests a higher likelihood of a token in that context, which allows the confidence of the LLM in its response to be gauged.
The response from the LLM was limited to either ‘Yes’ or ‘No’, and this limitation was put in place so that the log probability of the response could provide meaningful data. With this limitation in place, two aspects of the response were guaranteed:
  1. (1)
@@ 4.6. Logprobs
  2. (2)
    The response will be either ‘Yes’ or ‘No’
Since there is no variation in token length, that confounding variable is controlled against. In fact, this limitation is conceived after running a previous iteration of the study that did not limit the token length of the response. The data gathered from this initial run is extremely inconsistent, as the log probability of every token is affected by its context.It proves impossible to determine which tokens are significant and which are superfluous, so the limitation to keep the response one token long is implemented.
Since the LLM could only respond with either Yes or No, and the correct answer is always ‘Yes’ - the base truth for all code snippets being guaranteed (see ‘Dataset’ section) - two things can be deduced from the response:
  1. (1)
@@
4.7. Selected LLMs System
@@
  • gpt-3.5-turbo-0613: A snapshot of the GPT-3.5 turbo model taken on June 13th 2023, representing a previous iteration of the GPT-3.5 turbo model.
  • gpt-3.5-turbo-0125: At the time of writing, the latest GPT-3.5 turbo model available. The results for this model demonstrate the impact of advancements made over a full year cycle.
  • gpt-4o-2024-05-13: The most advanced model available in the OpenAI API (gpt, 2024). This model represents the cutting edge of large language models, and conclusions can be drawn as to the different way it handles perturbation when compared to the older 3.5 generation models.
4.8. Dataset
@@
To identify a large-scale dataset suitable for this study, inspiration is taken from a similar study (Jalil et al., 2023) which also aimed to prompt ChatGPT with software-related questions. In that study, ChatGPT is tasked with answering questions from a ‘popular software testing textbook’. Since this study needs a greater number of code snippets, that particular solution is not followed for implementation, but a similar solution is arrived at.
@@
A dataset of 2644 LeetCode coding questions is found through huggingface.co (hug, 2023), an AI community website featuring user-submitted models and datasets. This dataset is deemed appropriate due to containing code snippets as well as their related descriptions. This means that it would not be necessary to find a ‘base truth’ for all code snippets due to the veracity of provided descriptions being effectively guaranteed by LeetCode. The questions in the dataset have three categories of difficulty: easy, medium, and hard. Each question has answers available in four different coding languages: java, c++, python, and JavaScript. For this experiment, only JavaScript code snippets are used, to control for the confounding variable of different languages.
5. Results
5.1. Overview
@@
Figure 2. Correctness vs. perturbation budget for the Code Deletions attack across three models.
@@
We investigated the robustness of three GPT-based language models, two GPT-3.5 variants and one GPT-4o version, against four code perturbation methods: reordering, invisible characters, deletions, and homoglyphs. Our findings show that all models consistently achieve 100% correctness on unaltered (clean) code, demonstrating their baseline ability to comprehend straightforward snippets. However, the GPT-3.5 models generally display a uniform, step-wise decline in confidence and correctness as the perturbation budget increases, particularly for deletions and reordering, while invisible characters and homoglyphs prove slightly less disruptive. In contrast, GPT-4o exhibits a drastic performance drop as soon as any level of perturbation is introduced, yielding near-zero correctness in most cases and indicating that the newer model may have additional filtering or checks that invalidate the perturbed code. These trends highlight how different perturbation types undermine model confidence at varying rates and underscore the stark contrast in failure modes between the GPT-3.5 and GPT-4o generations.
5.2. GPT-3.5-turbo-0613
@@
From the Table 2, we can immediately find many similarities across the 4 categories. As expected, the confidence score for the clean code snippet is high with none below 95%, and the correctness is 100% across the board. The highest confidence score is held by invisible characters at 96.11%. Each category has their largest drop in confidence between clean and 20% perturbation budget, with differences getting smaller as the perturbation budget rises. The largest difference between budget levels is held by the deletions category between clean and 20%, at -52.76%. Reorderings and deletions have significantly lower floors for both confidence and correctness scores compared to invisible characters and homoglyphs, both approaching 0% for confidence, and both sub-20% for correctness. Deletions holds the lowest score across all perturbation budget levels for both confidence at 1.74% as well as correctness at 15.3%. The highest scores at 100% perturbation budget are held by the homoglyph category, with confidence at 54.23% and correctness at 67.67%. The homoglyph category is also the only perturbation category to not show decreases in scores at every level, with an increase in both confidence and correctness scores recorded between 60%-100%.
@@
Table 2. Performance of the model GPT-3.5-turbo-0613 under four imperceptible coding character attack methods.
5.3. GPT-3.5-turbo-0125
@@
Table 3. Performance of the model GPT-3.5-turbo-0125 under four imperceptible coding character attack methods.
5.4. GPT-4o-2024-05-13
Table 4 displays a clear and significant difference to the two GPT-3.5 models. However, there is one characteristic held in common across the results of all three models. Across all 4 categories, the clean code is still 100% correct and still holds the highest confidence score, with all of them holding a score of 81±1%. However, from there, things change drastically. The GPT-4o results are the only results to show negative confidence scores, which is due to the disproportionate number of incorrect answers, with the lowest score of -19.08% held by reorders at 20% perturbation budget. As with confidence scores, correctness is also extremely low across the board, with the next highest score following 100% being 2.33%, also held by reorders at 100% perturbation budget. The lowest correctness score is tied between reorders at 20% perturbation budget and invisible characters at 100% perturbation budget, and is also the lowest possible correctness, 0%.
@@ 5.4. GPT-4o-2024-05-13
Table 4. Performance of the model GPT-4o-2024-05-13 under four imperceptible coding character attack methods.
5.5. Answer to research questions
@@
To answer this question, we take the term ‘performance outcomes’ to refer to the correctness of a model’s responses. The introduction of perturbations can be found to greatly affect the correctness of the model responses, which can be gleaned simply from looking at the tables containing the summary of results. Statistical analysis confirms this conclusion using Pearson’s correlation coefficient, with a moderate-to-strong negative correlation coefficient found between perturbation budget and response correctness for all perturbation categories, as shown in Table. 5. Simply put, this means that as perturbation goes up, correctness goes down proportionally. The two GPT-3.5 generation models have very similar results, with both appearing to exhibit significantly higher negative correlation than the GPT-4o model. If only considering the correlation coefficient, it seems as if the GPT-4o model also shows a very high degree of normalization between the four categories of perturbation compared to the two 3.5 models. However, knowing that the results show a distribution discrepancy between the GPT-3.5 generation models and the GPT-4o model, we can investigate further.
@@
Table 5. Correlation coefficient of correctness scores against perturbation budget
If we consider the entire range of perturbation budget levels – i.e.: 0% - 100% – the correlation coefficient for the GPT-4o model results stands at around -0.6 for all categories – a moderately strong negative correlation. However, if we discard the results for the clean code, and only consider the perturbation budget levels greater than 0% – i.e.: we only consider code that has been perturbed – we find strong positive correlation for the reordering and deletion categories, whilst maintaining strong negative correlation for the remaining categories.
Clean Reord. Invis. Del. Homog.
Correct/1500 1500 12 15 14 10
Table 6. GPT-4.0 proportion of correct responses to incorrect responses
Table. 6 shows the proportion of total correct responses to incorrect responses returned by the GPT-4o model across all perturbation budgets. Since each category of perturbation creates five different levels of perturbation, and this study works with a 300-length dataset, each category of perturbation will make 1500 total prompts. Additionally, since the clean code is prompted with every perturbed code snippet, rather than only once per code-description pair, the clean prompt is made 1500 times too. This graph reveals that for the GPT-4o model responses, any correct response is in fact an outlier as the number of correct responses does not represent a significant proportion of the total number of responses, and it is the reason for the contradictory correlation coefficient exhibited by the statistical analysis.
Clean Reord. Invis. Del. Homog.
Correct/1500 1500 656 819 510 1064
Table 7. GPT-3.5-turbo-0613 proportion of correct responses to incorrect responses
Clean Reord. Invis. Del. Homog.
Correct/1500 1500 496 866 287 1038
Table 8. GPT-3.5-turbo-0125 proportion of correct responses to incorrect responses
For reference, Tables. 7 and 8 show the proportion of total correct responses to incorrect responses returned by the two GPT-3.5 models. It is immediately obvious that for these models, there is a statistically significant number of correct responses for each category.
This information also explains why the correlation coefficient for GPT-4o is so similar across all four categories of perturbation when the control is considered, whilst the GPT-3.5 models exhibit a much higher degree of variance. However, despite these limitations in GPT-4o results, statistical analysis can still conclusively demonstrate that the performance outcomes of the model are affected by perturbation, even if we cannot find a linear correlation between the correctness scores against the perturbation budget levels.
RQ2: Does a correlation exist between imperceptible perturbations and LLMs’ ‘confidence’, and what implications does this have for software testing?
@@
Table 9. Correlation coefficient of confidence scores against perturbation budget
Similarly to the correctness scores, the results exhibits shows a clear negative correlation between confidence scores and perturbation, which is again corroborated by statistical analysis using Pearson’s correlation coefficient. Likewise, there is a clear discrepancy between the results of the GPT-3.5 generation models and the GPT-4o model when clean responses are discarded, with the underlying reason being the same as for RQ1.
@@
Figure 3. Average difference in confidence scores for all models
@@
As with the previous research questions, the GPT-4o model contradictory results. All categories show close to -100% average difference for both overall confidence and overall correctness scores. This means that the overwhelming majority of responses were incorrect – previously established by Table 6. We can also conclude that GPT-4o does not differentiate between categories of perturbation, nor perturbation budgets – the results are the same whether 20% perturbed or 100% perturbed.
@@
However, even though the response is classified as ‘incorrect’ by the criteria of the study, it does not denote that the incorrect response is an undesirable outcome. A naïve conclusion drawn from the results of this study would be that the older generation of models are simply better at comprehending perturbed code than the newest generation, but the limitations of the study design mean that we cannot determine for what reason the GPT-4o model returned a ‘No’ answer. The more useful conclusion we should draw from the results is that the GPT-4o model can detect – with very high accuracy – the presence of special characters, recognizes their potential to confound the performance outcomes, and then deterministically returns a response that guarantees no false positives will be presented to the user.
6. Discussion
6.1. Novel insights
The purpose of this study is to contribute to the depth of research in large language model perturbation, rather than the breadth. Critical insight is gained into the inner mechanisms of three iterations of the most popular large language model in use (Karko, 2023), which has been widely adopted in certain security-sensitive sectors whilst being closed-source.
The process of undertaking this study has also provided insight into how the experiment methodologies of subsequent studies should be designed. Knowing now that the ‘flagship’ OpenAI product (gpt, 2024) exhibits behaviour that indicates the existence of deterministic guardrails put in place between the 3.5 and 4 generation models, we hope that ensuing research can investigate the finer details of such guardrails, and their effects on a system that is designed to be interacted with using natural language. Guardrails alone can negate all false positives, but a truly sophisticated system should recognize the presence of confounding characters, parse the code according to the spirit of the prompt, and return a correct answer, nonetheless; there should be no correlation between perturbation and correctness at all. However, in the absence of such a sophisticated system, it is encouraging to see that the evolution from 3.5-generation GPT to 4-generation GPT brings with it an unequivocally more consistent model.
6.2. Contributions and implications

### v3

@@ III-D Attack generation
Algorithm 1 formalises this perturbation process across four perturbations, providing detailed design to generate adversarial examples for downstream evaluation. Further details can be found in the replication package.
Algorithm 1 Imperceptible Code Perturbation Process
0: Clean code $C$ (length $L$), perturbation type $T\in\{\texttt{REORDER},\texttt{INVISIBLE},\texttt{DELETE},\texttt{HOMOGLYPH}\}$, perturbation budget $b\in[0,1]$
0: Perturbed code $C^{\prime}$
[…]
3:if $T=\texttt{REORDER}$ then
4:  prefix $\leftarrow\text{reverse}(C[0{:}n])$
5:  suffix $\leftarrow C[n:]$
6:  $C^{\prime}\leftarrow\text{unicodeReverse}(\text{prefix})+\text{suffix}$
7:  return $C^{\prime}$
8:end if
[…]
10:for $i=0$ to L1L-1 do
11:  $c\leftarrow C[i]$
12:  if $T=\texttt{INVISIBLE}$ and $count<n$ then
13:   $C^{\prime}\leftarrow C^{\prime}+c+\texttt{U+200C}$
14:  else if $T=\texttt{DELETE}$ and $count<n$ then
15:   $C^{\prime}\leftarrow C^{\prime}+\texttt{"a"}+\texttt{U+0008}+c$
16:  else if $T=\texttt{HOMOGLYPH}$ and $count<n$ then
17:   $C^{\prime}\leftarrow C^{\prime}+\text{homoglyph}(c)$
18:  else
19:   $C^{\prime}\leftarrow C^{\prime}+c$
20:  end if
21:  $count\leftarrow count+1$
22:end for
23:return $C^{\prime}$
III-E Prompting
We craft the prompts with the zero-shot prompting technique. We ensure that the templates of each prompt remain consistent across all categories of perturbations and all models. Following the implementation of different attack strategies, we leverage the Chat Completions API provided by the OpenAI to access the tested LLMs, which can be found in the replication package. The prompt template consists of the following three messages:
  1. 1.
@@ III-E Prompting
We note a similar question was explored in the study of [Fang24], which is open-ended in nature. However, our methodology adopts a more constrained approach by rephrasing the question to elicit a one-word response. Moreover, both the code and accompanying descriptions in our study are sourced from the LeetCode dataset [huggingface]. As such, we treat the description as the “ground truth”, given the curated process of the dataset. The perturbations are introduced to the code that preserve its overall semantics from a human perspective. This experiment design ensures that the correct answer to the question is consistently “Yes”, allowing for a controlled and reliable evaluation of LLMs’ comprehension.
III-F Logprobs
To present a thorough data analysis, we used the “logprobs” parameter of the Chat Completions API as the main vector. With logprobs enabled, the Chat Completions API returns the log probabilities of each output token in the response. Log probabilities of output tokens indicate the likelihood of each token occurring in the sequence [logprobs]. A higher log probability suggests a higher likelihood of a token in that context, which allows the confidence of the LLM in its response to be gauged.
To ensure meaningful and fair interpretation of these values, we have constrained the model’s output to a single token, which is either “Yes” or “No”. It guarantees that the log probability of the response could provide meaningful data. This experimental design provides two aspects of the response:
  1. 1.
@@ III-F Logprobs
  2. 2.
    The response will be either “Yes” or “No”
Since there is no variation in token length, our experiments control for the confounding effects introduced by multi-token outputs. In fact, this constraint is motivated by observations from a preliminary study of the experiment, in which unconstrained responses resulted in inconsistent and noisy logprobs results due to contextual dependencies across multiple tokens. It proves impossible to distinguish meaningful tokens from superfluous ones, prompting the adoption of this one-token constraint.
Given that the LLM must respond with either “Yes” or “No”, and the correct answer is always “Yes” which is guaranteed by the curated dataset, the response enable two key measurements:
  1. 1.
@@
III-G Confidence Score Calculation
To evaluate model confidence, we utilise the log probabilities returned by the OpenAI Chat Completions API with the logprobs=true setting. For each response, we convert the log-scale values into linear probabilities using the exponential function, average them across all tokens, and scale the result to a percentage range. This yields a normalised confidence score that reflects the model’s average token-level certainty in generating its full response.
Let $\log P_{i}$ be the log probability of the $i$-th token in a response, and let $N$ be the total number of tokens. The confidence for a given response is computed as:
$\text{Confidence}=\left(\frac{1}{N}\sum_{i=1}^{N}e^{\log P_{i}}\right)\times 100$
To measure the degradation caused by perturbations, we center each perturbed response’s confidence around the average confidence of clean (unperturbed) responses. For a perturbed response, we define the sample-level adjusted score as:
$\text{Score}_{\text{perturb}}=\begin{cases}\text{Conf}_{\text{perturb}}-\text{AvgCleanConf},&\text{if correct}\\ -100,&\text{if incorrect}\end{cases}$
This scheme penalises incorrect predictions sharply, ensuring they strongly influence the average. For correct responses, the score reflects relative changes in confidence without distorting scale.
Finally, to report confidence under each perturbation level in the results tables, we re-center the aggregate score by adding back the average clean confidence:
$\text{DisplayedConf}_{\text{perturb}}=\text{AvgCleanConf}+\frac{1}{N}\sum_{i=1}^{N}\text{Score}_{\text{perturb}}^{(i)}$
This formulation ensures that the reported confidence remains on the same scale as clean outputs, while still reflecting both prediction accuracy and confidence degradation under perturbation. Confidently incorrect answers are heavily penalised, while correct but uncertain answers contribute proportionally.
III-H Selected LLMs System
@@
  • gpt-3.5-turbo-0613: A snapshot of the GPT-3.5 turbo model released in June 2023, used to represent an earlier iteration of the GPT-3.5 generation.
  • gpt-3.5-turbo-0125: A more recent variant of GPT-3.5 turbo released in January 2024, included to reflect incremental improvements within the same generation.
  • gpt-4o-2024-05-13: A GPT-4 series model representing a newer generation with enhanced performance characteristics [gpt4o]. This allows for comparative analysis between 3.5-series models and a more advanced architecture in handling code perturbations.
III-I Dataset
@@
To identify a large-scale dataset suitable for this study, inspiration is taken from a similar study [Jalil23] which prompted ChatGPT with software-related questions. In that study, ChatGPT is tasked with answering questions from a “popular software testing textbook”. While this approach offers insightful findings, it is limited in scale for textbook questions. Differently, we adopt a similar solution to a broader and more scalable dataset, particularly with a greater number of code snippets.
@@
Herein, we leverage the LeetCode dataset from huggingface.co [huggingface], which is an AI community website featuring user-submitted models and datasets. It provides a rich set of code-description pairs for large-scale evaluation. In this work, we utilize a dataset of 2,644 LeetCode coding questions. Given the curated nature of LeetCode, the code-description pairs are assumed to be reliable and trustworthy, without requiring a “base truth” for code snippets. The questions in the dataset have three categories of difficulty: easy, medium, and hard. Each question has answers available in four different coding languages: Java, C++, Python, and JavaScript. For this experiment, only JavaScript code snippets are used, to control for the confounding variable of different languages.
IV Results
@@
Figure 2: Correctness vs. perturbation budget for the Code Deletions attack across three models.
IV-A Overview
@@
We investigated the robustness of three GPT-based language models, two GPT-3.5 turbo variants (GPT-3.5-turbo-0613 and GPT-3.5-turbo-0125) and one GPT-4o version (GPT-4o-2024-05-13), against four code perturbation methods: reordering, invisible characters, deletions, and homoglyphs. Our findings show that all models consistently achieve 100% correctness on unaltered (clean) code, demonstrating their baseline ability to comprehend straightforward snippets. However, the GPT-3.5 models generally display a uniform, step-wise decline in confidence and correctness as the perturbation budget increases, particularly for deletions and reordering, while invisible characters and homoglyphs prove slightly less disruptive. In contrast, GPT-4o exhibits a drastic performance drop as soon as any level of perturbation is introduced, yielding near-zero correctness in most cases and indicating that the newer model may have additional filtering or checks that invalidate the perturbed code. These trends highlight how different perturbation types undermine model confidence at varying rates and underscore the stark contrast in failure modes between the GPT-3.5 and GPT-4o generations.
IV-B Results on GPT-3.5-turbo-0613
@@
From the Table II, we can immediately find many similarities across the 4 categories. As expected, the confidence score for the clean code snippet is high with none below 95%, and the correctness is 100% across the board. The highest confidence score is held by invisible characters at 96.11%. Each category has their largest drop in confidence between clean and 20% perturbation budget, with differences getting smaller as the perturbation budget rises. The largest difference between budget levels is held by the deletions category between clean and 20%, at -52.76%. Reorderings and deletions have significantly lower floors for both confidence and correctness scores compared to invisible characters and homoglyphs, both approaching 0% for confidence, and both sub-20% for correctness. Deletions holds the lowest score across all perturbation budget levels for both confidence at 1.74% as well as correctness at 15.3%. The highest scores at 100% perturbation budget are held by the homoglyph category, with confidence at 54.23% and correctness at 67.67%. The homoglyph category is also the only perturbation category to not show decreases in scores at every level, with an increase in both confidence and correctness scores recorded between 60%-100%.
TABLE II: Performance of the model GPT-3.5-turbo-0613 under four imperceptible coding character attack methods.
@@
IV-C Results on GPT-3.5-turbo-0125
@@
TABLE III: Performance of the model GPT-3.5-turbo-0125 under four imperceptible coding character attack methods.
@@
IV-D Results on GPT-4o
Table IV displays a clear and significant difference to the two GPT-3.5 turbo models. However, there is one characteristic held in common across the results of all three models. Across all 4 categories, the clean code is still 100% correct and still holds the highest confidence score, with all of them holding a score of 81±1%. However, from there, things change drastically. The GPT-4o results are the only results to show negative confidence scores, which is due to the disproportionate number of incorrect answers, with the lowest score of -19.08% held by reorders at 20% perturbation budget. As with confidence scores, correctness is also extremely low across the board, with the next highest score following 100% being 2.33%, also held by reorders at 100% perturbation budget. The lowest correctness score is tied between reorders at 20% perturbation budget and invisible characters at 100% perturbation budget, and is also the lowest possible correctness, 0%.
TABLE IV: Performance of the model GPT-4o-2024-05-13 under four imperceptible coding character attack methods.
@@ IV-D Results on GPT-4o
IV-E Answer to research questions
@@
To answer this question, we take the term “performance outcomes” to refer to the correctness of a model’s responses. The introduction of perturbations can be found to greatly affect the correctness of the model responses, which can be gleaned simply from looking at the tables containing the summary of results. Statistical analysis confirms this conclusion using Pearson’s correlation coefficient, with a moderate-to-strong negative correlation coefficient found between perturbation budget and response correctness for all perturbation categories, as shown in Table V. Simply put, this means that as perturbation goes up, correctness goes down proportionally. The two GPT-3.5 generation models have very similar results, with both appearing to exhibit significantly higher negative correlation than the GPT-4o model. If only considering the correlation coefficient, it seems as if the GPT-4o model also shows a very high degree of normalization between the four categories of perturbation compared to the two 3.5 models. However, knowing that the results show a distribution discrepancy between the GPT-3.5 generation models and the GPT-4o model, we can investigate further.
TABLE V: Correlation coefficient of correctness scores against perturbation budget
@@
If we consider the full range of perturbation budgets, from 0% to 100%, the GPT-4o model results exhibit a correlation coefficient of approximately -0.6 across all categories, indicating a moderately strong negative correlation. However, when we exclude the clean code and focus only on cases where the perturbation budget is greater than 0%, we observe a strong positive correlation in the reordering and deletion categories, while the remaining categories continue to show a strong negative correlation.
TABLE VI: Number of correct responses (out of 1500) across different GPT models and perturbation types
Model Clean Reord. Invis. Del. Homog.
GPT-4o 1500 12 15 14 10
GPT-3.5-turbo-0613 1500 656 819 510 1064
GPT-3.5-turbo-0125 1500 496 866 287 1038
Table VI shows the proportion of total correct responses to incorrect responses returned by the GPT-4o model across all perturbation budgets. Since each category of perturbation creates five different levels of perturbation, and this study works with a 300-length dataset, each category of perturbation will make 1500 total prompts. Additionally, since the clean code is prompted with every perturbed code snippet, rather than only once per code-description pair, the clean prompt is made 1500 times too. The results reveal that for the GPT-4o model responses, any correct response is in fact an outlier as the number of correct responses does not represent a significant proportion of the total number of responses, and it is the reason for the contradictory correlation coefficient exhibited by the statistical analysis.
For reference, Table VI also shows the proportion of total correct responses to incorrect responses returned by the two GPT-3.5 models. It is immediately obvious that for these models, there is a statistically significant number of correct responses for each category. This information also explains why the correlation coefficient for GPT-4o is so similar across all four categories of perturbation when the control is considered, whilst the GPT-3.5 models exhibit a much higher degree of variance. However, despite these limitations in GPT-4o results, statistical analysis can still conclusively demonstrate that the performance outcomes of the model are affected by perturbation, even if we cannot find a linear correlation between the correctness scores against the perturbation budget levels.
RQ2: Does a correlation exist between imperceptible perturbations and LLMs’ “confidence”, and what implications does this have for software testing?
Similarly to the correctness scores, the results from Table VII exhibits shows a clear negative correlation between confidence scores and perturbation, which is again corroborated by statistical analysis using Pearson’s correlation coefficient. Likewise, there is a clear discrepancy between the results of the GPT-3.5 generation models and the GPT-4o model when clean responses are discarded, with the underlying reason being the same as for RQ1.
TABLE VII: Correlation coefficient of confidence scores against perturbation budget
@@
Figure 3: Average difference in confidence scores for all models
@@
As with the previous research questions, the GPT-4o model contradictory results. All categories show close to -100% average difference for both overall confidence and overall correctness scores. This means that the overwhelming majority of responses were incorrect, previously established by Table VI. We can also conclude that GPT-4o does not differentiate between categories of perturbation, nor perturbation budgets. And the results are the same whether 20% perturbed or 100% perturbed.
@@
However, even though the response is classified as “incorrect” by the criteria of the study, it does not denote that the incorrect response is an undesirable outcome. A naïve conclusion drawn from the results of this study would be that the older generation of models are simply better at comprehending perturbed code than the newest generation, but the limitations of the study design mean that we cannot determine for what reason the GPT-4o model returned a “No” answer. The more useful conclusion we should draw from the results is that the GPT-4o model can detect (with very high accuracy) the presence of special characters, recognises their potential to confound the performance outcomes, and then deterministically returns a response that guarantees no false positives will be presented to the user.
V Discussion
V-A Novel insights
The purpose of this study is to contribute to the depth of research in large language model perturbation, rather than the breadth. Critical insight is gained into the inner mechanisms of three iterations of the most popular large language model in use, which has been widely adopted in certain security-sensitive sectors whilst being closed-source.
The process of undertaking this study has also provided insight into how the experiment methodologies of subsequent studies should be designed. Knowing now that the “flagship” OpenAI product [gpt4o] exhibits behaviour that indicates the existence of deterministic guardrails put in place between the 3.5 and 4 generation models, we hope that ensuing research can investigate the finer details of such guardrails, and their effects on a system that is designed to be interacted with using natural language. Guardrails alone can negate all false positives, but a truly sophisticated system should recognise the presence of confounding characters, parse the code according to the spirit of the prompt, and return a correct answer, nonetheless; there should be no correlation between perturbation and correctness at all. However, in the absence of such a sophisticated system, it is encouraging to see that the evolution from 3.5-generation GPT to 4-generation GPT brings with it an unequivocally more consistent model.
V-B Contributions and implications